OpenStudy (anonymous):

"All three side lengths of the triangle are integers and together form a Pythagorean triple. Find the length of the third side and tell whether it is a leg or the hypotenuse" Question: 40 & 41

OpenStudy (anonymous):

So? What length did you find?

OpenStudy (anonymous):

40 and 41 are the lengths. And I need to find the third one.

OpenStudy (anonymous):

so 40^2+41^2=c^2

OpenStudy (anonymous):

But I'm getting a decimal.

OpenStudy (anonymous):

It didn't say if 41 or 40 was the hypotenuse or not. Try looking around for it. For example: 41^2=a^2+40^2. Play around with 41 and 40 being a, b, or c in the pythagorean theorem.

OpenStudy (anonymous):

It also said to figure out which length is the hypotenuse, which I don't know as well.

OpenStudy (anonymous):

Pythagorean Theorem solves for the hypotenuse given two legs. Just give it a try, if you can't do it I will try it with you.

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

remember that the largest value is always the hypotenuse.

OpenStudy (anonymous):

So if there were 2 numbers that had similar values, such as 40 and 41, could it be certain that they are both the legs?

OpenStudy (anonymous):

Not unless you knew the third value. The hypotenuse could be something like 57.28.

OpenStudy (anonymous):

Okay. Seems reasonable.

OpenStudy (triciaal):

41^2 = 1681 40^2 = 1600 difference of 81 = 9^2 as stated above the hypothenuse is the longest side, the largest value